OpenStudy (anonymous):

Solve the system of linear equations graphically, and name the type of system. 4x - y = -4 12x - 3y =24

OpenStudy (anonymous):

Solve for y for both equations. The graph it out, and the point of intersection is your answer...I think.

OpenStudy (kropot72):

4x - y = -4 -y = -4x - 4 y = 4x + 4 ....................(1) 12x - 3y = 24 -3y = -12x + 24 y = 4x - 8 ....................(2) When the two equations are rearranged as above it is clear that there is no solution. The slopes are equal so the lines never intersect.